# Motion-Aware Generative Frame Interpolation

### v1

@@ Abstract
Generative frame interpolation, empowered by large-scale pre-trained video generation models, has demonstrated remarkable advantages in complex scenes. However, existing methods heavily rely on the generative model to independently infer the correspondences between input frames—an ability that is inadequately developed during pre-training. In this work, we propose a novel framework, termed Motion-aware Generative frame interpolation (MoG), to significantly enhance the model’s motion awareness by integrating explicit motion guidance. Specifically we investigate two key questions: what can serve as an effective motion guidance, and how we can seamlessly embed this guidance into the generative model. For the first question, we reveal that the intermediate flow from flow-based interpolation models could efficiently provide task-oriented motion guidance. Regarding the second, we first obtain guidance-based representations of intermediate frames by warping input frames’ representations using guidance, and then integrate them into the model at both latent and feature levels. To demonstrate the versatility of our method, we train MoG on both real-world and animation datasets. Comprehensive evaluations show that our MoG significantly outperforms the existing methods in both domains, achieving superior video quality and improved fidelity.
Figure 1: Examples of frame interpolation in real-world and animation scenes. Compared to other methods, our approach, MoG, exhibits superior stability in motion and consistency in content relative to the input frames. For optimal viewing, please zoom in.
@@ 1 Introduction
Video Frame Interpolation (VFI), which seeks to synthesize intermediate frames between two input frames, has garnered significant attention in recent years due to its capacity for enhancing video frame rates in video post-processing. Flow-based VFI methods [10, 41, 42, 15, 13] predominantly rely on correspondence information through estimating the motion between input frames—termed intermediate flow—to warp information and generate intermediate frames. However, these approaches are primarily optimized for rigid motion scenarios. Consequently, when confronted with complex scenes, the interpolated frames often exhibit pronounced blurring and artifacts, as the results of the flow-based method EMA-VFI [41] shown in Fig. 1.
@@
To overcome these limitations, recent advancements [34, 35, 33, 3] have shifted towards leveraging video generation models [35, 2] for frame interpolation, capitalizing on their strong generative capabilities in dynamic scenes. Despite the notable improvements exhibited in complex scenarios [34], current approaches still face a key challenge: they fail to explicitly exploit the dynamics between the two input frames. Instead, they solely rely on the generative model to infer the correspondences between input frames by itself—an ability that is insufficiently nurtured during the generative pre-training. Consequently, the videos generated by these methods often suffer from unstable motion and incoherent content to input frames, as shown by DynamiCrafter [35] and ToonCrafter [34] in Fig. 1.
In this work, we introduce a new framework, Motion-aware Generative frame interpolation (MoG), marking the first explicit incorporation of motion guidance between input frames to enhance the motion awareness of generative models. Our approach significantly alleviates the difficulty of inferring the dynamics between input frames to generate the realistic motion. To achieve this goal, we address two pivotal questions: what can serve as an effective motion guidance, and how to seamlessly and harmlessly integrate this guidance into the generative model.
For the first question, we propose leveraging the intermediate flow from flow-based frame interpolation models [41] as explicit motion guidance. The intermediate flow can be utilized to obtain coarse approximations of intermediate frames under the assumption of rigid motion, which can serve as effective guidance to inspire and stabilize the inference of realistic motion in the generation process. Moreover, compared to direct estimation by pre-trained optical flow models [36, 28, 37], the intermediate flow is task-oriented trained [38], making it inherently more suitable for generating intermediate frames. Flow-based interpolation models also specifically account for occlusion masks [41] between the two input frames, enabling a more accurate aggregation of information into the intermediate frame—something that optical flow often fails to achieve.
To address the second question, we propose a simple yet effective strategy: warping the information from input frames to derive the guidance-based intermediate frame representations, which are subsequently integrated into the pre-trained model as explicit motion cues. This integration is performed at both the latent and feature levels, enabling the model to be fully aware of the correspondences between input frames. Specifically, at the latent level, we concatenate a collection of auxiliary latents to the input at each denoising step [8], which includes the latents from the input frames as well as the warped intermediate latents. In terms of feature-level integration, we reuse the pre-trained temporal layers [35] to temporally smooth the warped features, ensuring their alignment within the original feature space, and then incorporate them into the network. Notably, our design requires no additional parameters and achieves improved performance with few fine-tuning steps.
To comprehensively evaluate the versatility of our method, we develop MoG for both the real-world and animation scenes. Furthermore, we meticulously curate two testing datasets, VFIBench-Real and VFIBench-Ani, to assess the performance of each scenes. Experimental results indicate that MoG significantly outperforms the existing generative interpolation models in both qualitative and quantitative aspects. Qualitatively, the videos generated by MoG exhibit enhanced motion stability and greater content consistency, as evidenced in Fig. 1. Quantitatively, MoG demonstrates the obvious improvements in video quality metrics and superior fidelity when compared to ground truth videos. Our contributions are summarized as follows:
  • We introduce a new generative frame interpolation framework, MoG, which is the first to incorporate motion guidance to enhance the motion awareness.
  • We demonstrate that the intermediate flow from flow-based VFI can effectively serve as a motion guidance, and we design a simple and efficient method to integrate this prior into the network at both the latent and feature levels.
@@
  • We validate the effectiveness of MoG in both real-world and animated scenarios, with experimental results showing that MoG significantly outperforms existing open-source generative frame interpolation methods.
@@ 2 Related Work
Figure 2: Overview of MoG. MoG consists of two parts. First, it extracts the motion guidance of the input frames using a pre-trained flow-based VFI model. Subsequently, this guidance is seamlessly injected into the generative model at both the latent and feature levels.
@@ 2.1 Flow-based Frame Interpolation
Flow-based video interpolation, explicitly estimating the intermediate flow from the input frames to the intermediate frame, have become dominant in deterministic frame interpolation [17]. It can be broadly categorized into two classes based on how the intermediate optical flow is derived. The first class [9, 19, 18, 20, 1, 12] utilized pre-trained optical flow models to obtain the intermediate flow either directly or through refinement. For instance, SoftSplat [19] linearly adjusted the bidirectional flow estimated by PWC-Net [27] to represent the intermediate flow and employs an improved forward warping to aggregate information. The second class [41, 10, 42, 15, 13, 14, 21, 16] modeled the correspondence information of the input frames to directly predict the intermediate flow, offering greater flexibility and task-oriented modeling capacity compared to the first approach. RIFE [10] demonstrated that simple convolutional layers can effectively predict the intermediate flow, achieving impressive efficiency. Similarly, EMA-VFI [41] enhanced the flow prediction by explicitly modeling the dynamics between frames through inter-frame cross-attention. However, both of classes are primarily designed under the assumption of rigid motion scenes. When confronted with complex motion scenarios, they often exhibit significant blurring and artifacts. In our work, we focus on the generative frame interpolation by leveraging the second class of flow-based video frame interpolation (EMA-VFI [41]) to provide explicit motion guidance.
@@ 2.2 Generative Frame Interpolation
Recent work has begun to explore the use of large-scale pre-trained video generation models [2, 35], which excel at generating videos in complex dynamic scenes, for the VFI task. Current generative frame interpolation methods can be categorized into two types: the first [3, 33] employed pre-trained generative models to perform image-to-video tasks conditioned on the initial and final frames, subsequently merging the resulting videos to create the final interpolated frame. For example, GI [33] enhanced the motion stability by controlling the consistency of temporal correlations across the two generation processes. The second category [35, 34, 15, 32, 44] focused on fine-tuning video generation models specifically for interpolation, by integrating information from input frames into the model’s architecture and optimizing it for end-to-end interpolation. DynamiCrafter [35] was trained for real-world interpolation, while ToonCrafter [34] was tailored for animated scenes. Although all these methods have demonstrated significant improvements in generating complex scenarios, they do not explicitly consider the correspondence between input frames, which complicates the motion inference of generative models. In contrast, we are the first to explicitly introduce the correspondence guidance to enhance the motion awareness of generative models and our method achieves superior video quality and fidelity.
@@ 3 Preliminaries
Figure 3: Comparison of different methods for approximating intermediate frames. Linear interpolation merely overlays the input frames, while optical flow often introduces significant artifacts. In contrast, the frames warped through intermediate flow demonstrate superior stability, providing the generative model with robust correspondence information.
@@ 3.3 VFI with Diffusion Models
The design of the denoising network $\epsilon_{\theta}$ follows an U-Net-like structure [26], consisting of contracting blocks, middle blocks, and expansive blocks. Each block comprises spatial and temporal layers. The spatial layers mainly consist of ResNet blocks [6] and Transformer blocks [31] with spatial attention, modeling spatial information within each frame, while the temporal layers are formed by Transformer blocks with temporal self-attention.
@@ 4 Motion-Aware Generative VFI
Currently, most generative interpolation methods [33, 3, 35, 34, 5] conduct frame interpolation by directly incorporating information from the input frames into the model’s input [5, 35] or decoder [35]. However, in these methods, the correspondence between input frames, which is crucial for understanding and reasoning about the motion and appearance of the intermediate frames, have to be inferred independently by the generative model itself. This correspondence estimation capability has not been adequately cultivated in the pre-training of large-scale text-to-video generation or pure video generation tasks. Consequently, this insufficient exploration of correspondence leads to inconsistencies in motion and appearance of the generated videos relative to the input frames.
To address this issue, we propose Motion-Aware Generative Frame Interpolation (MoG), the first approach to explicitly introduce correspondence information between the two frames into the generative model, thereby reducing the difficulty of the inference for motion and appearance for generative model. As illustrated in Fig. 2, MoG comprises two parts: it first extracts an explicit motion guidance based on the input frames by a pre-trained flow-based VFI model [41]; it then seamlessly injects this guidance into the generative model at both the latent and feature levels. Notably, our method does not require any additional parameters; it merely necessitates minimal fine-tuning steps on spatial layers of the denoising network to compensate for changes in feature distribution.
4.1 Choice of motion guidance
As illustrated in Fig. 3, in addition to the intermediate flow described in Eq. 1, there are two other guidance for representing intermediate frames based on input frames. The first method is linear interpolation, which directly weights the representations of the initial and final frames according to the time step of the intermediate frame. While straightforward, this technique merely overlays the input frames, offering limited correspondence information. The second method employs a pre-trained optical flow estimator, such as [37], to derive the flow between input frames, which is then scaled to generate the intermediate flow. This approach often suffers from limited generalization due to being predominantly trained on synthetic datasets. Moreover, it cannot account for occlusion relationships between input frames. This can lead to artifacts in complex scenes, such as jagged edges and ghosting effects.
In contrast, the intermediate flow defined in Eq. 1 is specifically trained for the VFI task, making it inherently more suitable for representing the correlation between intermediate frames and the initial and final frames. It also predicts an occlusion mask $M^{n}$ for each frame. As illustrated in Fig. 3, the intermediate flow allows for the approximation of intermediate frames, which, despite some blurriness, maintains stable motion trajectories and consistent object appearances. This stability alleviates the difficulties of real motion inference in generative models. This is further validated by our experiments in Sec. 5.4.
4.2 Motion Guidance Integration
Due to the inability of diffusion models to directly use correspondence guidance, it is necessary to devise a strategy to inject these motion informaiton into the denoising network. To achieve this, we propose a simple yet effective approach that leverages the guidance, akin to the Eq. 2, to coarsely estimate the representation of the intermediate frames from the representations of the start and end frames. This estimated representation is then seamlessly merged into the denoising network. Furthermore, to fully exploit the correspondence guidance across different granularities of the generation process, we conduct the guidance injection at both the latent and feature levels.
@@ 4.2 Motion Guidance Integration
To introduce motion guidance at the latent level, we propose to coarsely estimate the latent code of intermediate frames by backward warping of the input frame’s latent code with the intermediate flow. Specifically, given the latent codes of the start and end frames $z_{0}^{1}$ and $z_{0}^{N}$, along with the motion guidance obtained via Eq. 1, we estimate the latent code of the $n$-th intermediate frame:
@@
It is noteworthy that no additional parameters are required, as our base models, DynamiCrafter or ToonCrafter, have already designed to accommodate extra inputs; however, the latent codes of intermediate frames in $\mathbf{\bar{z}}_{0}$ are always zero at their methods.
@@
To effectively integrate the motion guidance in different granularities, we propose to inject guidance also in feature-level. Analogous to the latent-level, we estimate the features of intermediate frames based on the features $F^{0},F^{N}\in\mathbb{R}^{D\times H\times W}$ of the input frames:
@@
In this equation, $\bar{F}^{i}$ represents the estimated features of the $i$-th frame under the correspondence prior. Unfortunately, unlike in latent-level injection, direct concatenation of the warped intermediate features into the network is not feasible. To address this issue, we leverage the temporal layer of the denoising network, which enables us to smooth and align the estimated intermediate features with the original feature distribution without introducing additional parameters, as shown in Fig. 2:
@@
Subsequently, we incorporate the smoothed features $\dot{F}$ into the original features:
$\mathbf{\tilde{F}}=(\mathbf{F}+\mathbf{\hat{F}})/2.$ (10)
Remarkably, our exploration (refer to Sec. 5.4) reveals that this simple averaging already allows the generative model to effectively utilize the introduced motion guidance.

### v2

@@ Abstract
Flow-based frame interpolation methods ensure motion stability through estimated intermediate flow but often introduce severe artifacts in complex motion regions. Recent generative approaches, boosted by large-scale pre-trained video generation models, show promise in handling intricate scenes. However, they frequently produce unstable motion and content inconsistencies due to the absence of explicit motion trajectory constraints. To address these challenges, we propose Motion-aware Generative frame interpolation (MoG) that synergizes intermediate flow guidance with generative capacities to enhance interpolation fidelity. Our key insight is to simultaneously enforce motion smoothness through flow constraints while adaptively correcting flow estimation errors through generative refinement. Specifically, we first introduce a dual guidance injection that propagates condition information using intermediate flow at both latent and feature levels, aligning the generated motion with flow-derived motion trajectories. Meanwhile, we implemented two critical designs, encoder-only guidance injection and selective parameter fine-tuning, which enable dynamic artifact correction in the complex motion regions. Extensive experiments on both real-world and animation benchmarks demonstrate that MoG outperforms state-of-the-art methods in terms of video quality and visual fidelity. Our work bridges the gap between flow-based stability and generative flexibility, offering a versatile solution for frame interpolation across diverse scenarios.
Figure 1: Examples of frame interpolation in real-world and animation scenes. Compared to other methods, our approach, MoG, exhibits superior stability in motion and consistency in appearance details.
@@ 1 Introduction
Video Frame Interpolation (VFI), which aims to synthesize intermediate frames between two input frames, has garnered significant attention in recent years due to its capacity for enhancing video frame rates in video post-processing. Flow-based VFI methods [9, 45, 46, 17, 14] predominantly rely on estimating the motion between input frames—termed intermediate flow—to warp condition information and generate intermediate frames. Conventionally, the generated intermediate frames follow the motion trajectories encoded in the estimated intermediate optical flow, yielding temporally coherent video sequences. However, in scenarios involving complex motions such as object deformations, accurate intermediate flow estimation becomes infeasible, leading to pronounced artifacts in corresponding regions. As demonstrated in Fig. 1 by the flow-based method EMA-VFI [45], while the overall results exhibit temporal smoothness, complex regions manifest blurring and ghosting artifacts.
@@
Recent advancements [38, 39, 37, 4] have increasingly focused on leveraging video generation models [39, 2] for frame interpolation, capitalizing on their robust generative capacities in dynamic scenes. While notable improvements have been demonstrated in complex scenarios [38], current approaches rely exclusively on generative models to infer inter-frame correspondences—a capability that remains underdeveloped during generative pre-training. Consequently, these methods often produce unstable motion patterns and inconsistencies with input frames due to the lack of explicit motion trajectory constraints, as illustrated by DynamiCrafter [39] and ToonCrafter [38] in Fig. 1.
In this work, we introduce a new framework, Motion-aware Generative frame interpolation (MoG), which integrates intermediate flow with generative capacities to to enhance interpolation fidelity. MoG is designed to to bridge the gap between flow-based stability and generative flexibility, offering a versatile solution for frame interpolation across diverse scenarios. Our core idea is to enforce motion smoothness via flow constraints while rectifying complex motion regions through generative refinement simultaneously. To attain this objective, we tackle two crucial questions: how to seamlessly incorporate flow constraints into the generative model, and how to endow the generative model with the ability to dynamically correct flow errors.
For the first question, to ensure that the generated motion trajectories are guided by the estimated intermediate flow, we introduce dual guidance injection. At both the latent and feature levels, we warp the information of the input frames using the intermediate flow and propagate it into the generation process of intermediate frames, serving as an explicit motion guidance for inferring motion. Notably, compared with ControlNet-like designs [47, 50], our design requires no additional parameters and thereby better preserves the pre-trained capabilities.
To address the second challenge, we introduce two critical designs: encoder-only guidance injection and selective parameter fine-tuning. Specifically, guidance is exclusively injected at the encoder stage of the generative model, enabling the decoder to adaptively adjust the generation process. Furthermore, we only fine-tune the spatial layers to adapt to guidance fusion while preserving motion modeling capabilities in temporal layers.
To thoroughly evaluate the versatility of our approach, we adapt MoG for both real-world and animation scenes. Experimental results demonstrate that MoG substantially outperforms existing generative interpolation models in both domains. Specifically, the interpolated videos generated by MoG exhibit superior motion stability and improved content consistency, as visually demonstrated in Fig. 1. Our contributions are summarized as follows:
  • We propose a novel frame interpolation framework, MoG, which is the first to bridge the gap between flow-based stability and generative flexibility.
  • We introduce dual-level guidance injection to constrain the generated motion with the motion trajectories derived from the flow.
  • We implement encoder-only guidance injection and selective parameter fine-tuning to endow the generative model with the ability to dynamically correct flow errors.
@@
  • Experimental results showing that MoG significantly outperforms existing generative frame interpolation methods in both qualitative and quantitative aspects.
@@ 2 Related Work
Figure 2: Overview of MoG. MoG consists of two parts. First, it extracts the intermediate flow between input frames. Subsequently, this guidance is seamlessly injected into the generative model at both the latent and feature levels. Meanwhile, the generative model would adaptively rectifying the errors by two crucial designs, namely, encoder-only guidance injection and selective parameter fine-tuning.
@@ 2.1 Flow-based Frame Interpolation
Flow-based video interpolation, explicitly estimating the intermediate flow from the input frames to intermediate frames, have become dominant in deterministic frame interpolation [19]. It can be broadly categorized into two classes based on how the intermediate optical flow is derived. The first class [8, 22, 21, 23, 1, 13] utilized pre-trained optical flow models to obtain the intermediate flow either directly or through refinement. For instance, SoftSplat [22] linearly adjusted the bidirectional flow estimated by PWC-Net [30] to represent the intermediate flow and employs an improved forward warping to aggregate information. The second class [45, 9, 46, 17, 14, 15, 24, 18] modeled the correspondence information of the input frames to directly predict the intermediate flow, offering greater flexibility and task-oriented modeling capacity compared to the first approach. RIFE [9] demonstrated that simple convolutional layers can effectively predict the intermediate flow, achieving impressive efficiency. Similarly, EMA-VFI [45] enhanced the flow prediction by explicitly modeling the dynamics between frames through inter-frame cross-attention. However, when confronted with complex motion scenarios, both of classes often exhibit significant blurring and ghosting artifacts. In this work, we leverage the intermediate flow to enhance the temporal smoothness of generative models. Meanwhile, we utilize generative models to rectify the errors of the intermediate flow in complex scenarios.
@@ 2.2 Generative Frame Interpolation
Recent work has begun to explore the use of large-scale pre-trained video generation models [2, 39], which excel at generating videos in complex dynamic scenes, for the VFI task. Current generative frame interpolation methods can be categorized into two types: the first [4, 37, 50, 43] employed pre-trained generative models to perform image-to-video tasks conditioned on the initial and final frames, subsequently merging the resulting videos to create the final interpolated frame. For example, GI [37] enhanced the motion stability by controlling the consistency of temporal correlations across the two generation processes. The second category [39, 38, 17, 35, 49, 3, 11, 31] focused on fine-tuning video generation models specifically for interpolation, by integrating information from input frames into the model’s architecture and optimizing it for end-to-end interpolation. DynamiCrafter [39] was trained for real-world interpolation, while ToonCrafter [38] was tailored for animated scenes. Although all these methods have demonstrated significant improvements in generating complex scenarios, they do not explicitly consider the correspondence between input frames, which complicates the motion inference of generative models. In contrast, we are the first to explicitly introduce the motion guidance to enhance the motion smoothness of generative models and our method achieves superior video quality and fidelity. Currently, there are also some works [42, 20, 16, 36] that use optical flow to assist generation in other tasks. They typically require the accurate optical flow between all adjacent frames to aid the generation process, whereas we only use the coarse intermediate flow between the two input frames to smooth the motion. Meanwhile, our method could dynamically correct the errors in the intermediate flow.
@@ 3 Preliminaries
Figure 3: Demonstration of MoG’s guidance correction capability. In both examples, DynamiCrafter or ToonCrafter struggle to generate temporally consistent motion in complex scenarios. While intermediate flow can provide valuable motion cues, it often introduces artifacts and fails to render fine appearance details. Leveraging the encoder-only injection design and selective parameter fine-tuning, MoG effectively integrates reliable motion information from intermediate flow while correcting its inaccuracies.
@@ 3.3 VFI with Diffusion Models
The design of the denoising network $\epsilon_{\theta}$ follows an U-Net-like structure [29], consisting of encoder blocks and decoder blocks. Each block comprises spatial and temporal layers. The spatial layers mainly consist of ResNet blocks [6] and Transformer blocks [34] with spatial attention, modeling spatial information within each frame, while the temporal layers are formed by Transformer blocks with temporal self-attention.
@@ 4 Motion-Aware Generative VFI
4.1 Motivation
Flow-based frame interpolation methods can generate temporally smooth videos with the assistance of intermediate flow. However, in regions with complex motions, accurate intermediate flow estimation remains challenging due to the lack of sufficient supervisory signals—the annotation of real-world optical flow is prohibitively expensive. This limitation leads to severe artifacts such as blurring and ghosting in flow-based approaches. In contrast, recent generative frame interpolation methods, benefiting from large-scale pre-training, demonstrate robust generative capabilities in complex scenarios. Unfortunately, they often suffer from incoherent motion and content inconsistencies with input frames. We attribute this to the fact that existing methods rely solely on generative models to infer motion trajectories between input frames—a capability insufficiently nurtured during generative pre-training. To bridge the gap between flow-based stability and generative flexibility, we introduce Motion-aware Generative frame interpolation (MoG), as illustrated in Fig. 2. MoG employs the intermediate flow from flow-based methods as an explicitly motion guidance for generation, smoothing the motion of generated videos. Concurrently, MoG leverages the refinement capacities of generative models to automatically correct flow errors in complex motion regions.
4.2 Dual Guidance Injection
Since generative models cannot directly utilize the intermediate flow, it is imperative to devise a strategy for sufficiently injecting motion guidance into the denoising network. To this end, we propose dual guidance injection. Similar to the operation in Eq. 2, we coarsely estimate the representation of the intermediate frames using the intermediate flow. Subsequently, the estimated representations are seamlessly integrated into the model at both the latent and feature levels.
@@ 4.2 Dual Guidance Injection
To incorporate flow constraints at the latent level, we propose to inject additional latent codes of intermediate frames into the input. We backward-warp the latent codes of the input frames, guided by the intermediate flow. Specifically, given the latent codes of the start and end frames $z_{0}^{1}$ and $z_{0}^{N}$, along with the motion guidance obtained through Eq. 1, we estimate the latent code of the $n$-th intermediate frame as follows:
@@
It is noteworthy that no additional parameters are required, as our base models, DynamiCrafter or ToonCrafter, have already designed to accommodate extra inputs; In contrast, in their methods, the latent codes of intermediate frames in $\mathbf{\bar{z}}_{0}$ are always set to zero.
@@
To effectively integrate the motion guidance cross different granularities, we propose to inject guidance also in feature-level. Analogous to the latent-level, we estimate the features of intermediate frames based on the features $F^{1},F^{N}\in\mathbb{R}^{D\times H\times W}$ of the input frames:
@@
In this equation, $\bar{F}^{i}$ represents the estimated features of the $i$-th frame guided by the intermediate flow. Unfortunately, unlike in latent-level injection, direct concatenation of the warped intermediate features into the network is not feasible. To address this issue, we reuse the temporal layer of the denoising network, which enables us to smooth and align the estimated intermediate features $\mathbf{\bar{F}}$ with the original feature distribution without introducing additional parameters. As shown in Fig. 2, the smoothed features $\mathbf{\hat{F}}$ is acquired by:
@@
Subsequently, we incorporate the flow-derived features $\dot{F}$ into the original features:
$\mathbf{\tilde{F}}=\phi(\mathbf{F},\mathbf{\hat{F}}).$ (10)
Remarkably, our exploration (refer to Sec. 5.4) reveals that a simple averaging already allows the generative model to effectively utilize the introduced motion guidance.
4.3 Guidance Correction
As mentioned in Sec. 4.1, the intermediate flow cannot be accurately predicted in regions with complex motions. To alleviate the potential adverse effects, we have implemented two crucial designs to endow the generative model with the ability to automatically rectify error-prone regions.
Firstly, we adopt the design of encoder-only guidance injection. As depicted in Fig. 2, we introduce feature-level guidance injection exclusively in the encoder blocks of the denoising network. In this way, the decoder can appropriately adjust and rectify the information from the encoder, thereby capitalizing more effectively on the valuable information embedded in the flow guidance. Secondly, we conduct selective parameter fine-tuning. We only fine-tune the spatial layers of all blocks. Given the inherent inaccuracy of the intermediate flow during the training phase, the model can learn how to dynamically harness the useful information within the guidance while concurrently acquiring the ability to repair flawed regions. Freezing the temporal layers serves to preserve the motion inference capability during pre-training and prevent the performance degradation associated with the feature-level guidance injection.
Discussion.
To validate the effectiveness of our proposed method, as shown in Fig. 3, we showcase two examples for a comparison among the results derived from the intermediate flow, those of the original generative model, and the results produced by MoG. Evidently, MoG can fully exploit the motion cues encapsulated within the intermediate flow to generate videos with better temporal smoothness. Meanwhile, in regions with complex motions, MoG rectifies the inaccuracies from the intermediate flow and leverages the generative capabilities to yield more reasonable appearance details. More quantitative comparisons can be found in Sec. 5.4.